# Deep Learning based Key Information Extraction from Business Documents: Systematic Literature Review

### v1

@@ 5.3.2. Quantitative comparison
We now discuss the evaluation results for the common benchmark datasets CORD, FUNSD and SROIE. We only consider the F1 score as it represents the most common evaluation metric. Note that we do not distinguish between micro, macro or weighted F1 averages [^7] For an explanation of these aggregated metrics, we refer to (Sokolova and Lapalme, 2009)., since in many cases it has not been explicitly specified by the authors. The results should therefore be treated with caution, especially since macro averages usually tend to be lower than the micro averages.
To this end, figure 6(c) shows the results for the three datasets given all papers that present results on the corresponding datasets. A trend line is also displayed. In all cases, one can see that there is no steep increase in performance over time. However, the largest improvement over time was achieved with FUNSD, as can be seen from the trend line. In the case of SROIE, the trend line is almost horizontal, therefore indicating a stagnation. All graphs show a few outliers that achieved worse results than previous and subsequent work. One reason for this may be a different calculation of the F1 scores, e.g., by determining macro average values instead of micro average ones. The best performing approach for CORD is #28 proposed by (Gu et al., 2021), for FUNSD #85 by (Li et al., 2023b) and regarding SROIE #96 by (Zhang et al., 2023c). The first two are sequence-based approaches, while the latter is a graph-based method. On average, very comparable results are obtained for CORD and SROIE, which can be attributed to the fact that they both contain receipts. The average F1 score for both datasets is 0.96. In case of FUNSD, however, the results are significantly worse, with an average F1 score of only 0.82. One problem that KIE approaches face with FUNSD may be that this dataset was not primarily constructed for the KIE task. The fields to be extracted often span multiple lines of text, which is very different from traditional KIE datasets that aim at extracting key information such as a specific date.
@@
Figure 6(c) visualizes the relation between model sizes in terms of parameter count and performance in terms of F1 score for the three datasets. Note the logarithmic scale of the x-axis due to the large range of possible values. In all cases, there is a clear cluster formation with some outliers. The correlation coefficient for CORD is -0,10 (p=0,62), for FUNSD 0,13 (p=0,49) and for SROIE 0,12 (p=0,65). Therefore, there is no significant correlation (neither negative nor positive) between model size and performance. Interestingly, the correlation in case of CORD is even negative, which means that many of the approaches with smaller parameter counts are able to outperform more complex KIE methods. The figure also differentiates the KIE paradigms by color. One interesting observation is that regarding FUNSD, especially graph-based methods often produce worse results compared to other paradigms and generally represent outliers. One reason for this could be the special layout of the forms in this dataset. As mentioned before, the forms typically contain entities that span multiple text lines. A constructed graph, independent of the chosen granularity, is indisputable an unfavorable document representation for such layouts. At the same time, one can see across all datasets (especially FUNSD), that a combination of graph-based and sequence-based modules seems to increase extraction performance compared to a purely graph-based approach. The figure indicate that increasing the model size beyond a certain threshold does not necessarily correlate with a significant increase in performance and that creating more complex models can even yield diminishing performance returns. The fact that all paradigms can achieve competitive results (with the exception of graph-based methods w.r.t FUNSD) also suggests that the choice of model architecture can be flexible depending on the specific application. Therefore, approach design seem to be more crucial than the sheer parameter count. This is a relevant insight especially for practical applications, where resources and computational cost need to be taken into account.
In case of FUNSD, the spread is much larger compared to the other two datasets. Here, many of the approaches with fewer parameters achieve significantly worse results compared to rest. One can conclude that the challenging FUNSD dataset with its unique entities therefore represents a particular challenge for KIE systems of lower complexity. The approach #81 proposed by (He et al., 2023b) with 175 billion parameters is not able to outperform most of the other approaches with significantly fewer parameters. This is especially true for CORD, where it is one of the worst performing approaches. However, competitive results are achieved for the other datasets, especially SROIE. Approach #75 by (Dhouib et al., 2023) with 70 million parameters represents another outlier and achieves less competitive results regarding CORD and SROIE, even though the model size is not significantly smaller compared to the other methods. The system by (Hamri et al., 2023) only contains 53,6 thousand parameters and is therefore the by far smallest one compared to the other analyzed KIE methods. Nonetheless, the approach can achieve convincing results in case of SROIE and even outperforms the aforementioned method with 70 million parameters.
@@ 6.1. Pre-Training and Fine-Tuning procedures
As discussed in chapter 5.2.1, the majority of authors adopt MVLM during pre-training. In some cases, it is even the only pre-training objective that is being used. Given the general popularity of MLM in the NLP domain, it is not surprising that the MVLM task, which is derived from it, is also popular in KIE research. It represents an effective method to let models learn useful representations given the surrounding textual and positional contexts. As the adequate processing of VRDs also requires to consider visual cues, newer pre-training objectives have been proposed in order to fuse visual information into KIE systems. This can be done in different ways, for example by letting the model reconstruct local or global document image snippets (Appalaraju et al., 2021; Tang et al., 2023) or predicting the segment length of an image snippet (Li et al., 2021c). Other possibilities aim at various matching-tasks between text and document images, for example where the model has to predict whether a sentence describes a document image (Appalaraju et al., 2021), if a given image and text are part of the same page (Xu et al., 2021) or if image patches of a word are masked or not (Huang et al., 2022). Some authors also define additional pre-training objectives to better exploit positional information, for example by the means of classification problems to estimate a tokens placement into a document image area (Li et al., 2021a; Wang et al., 2022b) or to estimate relative positional directions of an element to its neighbors (Li et al., 2021c). Research is also conducted related to a better understanding of numerical values and their relationships in business documents (Douzon et al., 2022). This can be beneficial for processing document types such as invoices, which usually contain various monetary values that are closely related to each other. To conclude, the proposed pre-training objectives are carefully designed to allow for an adequate fusion of textual, positional and visual inputs, which helps corresponding models to obtain a document understanding which can then be exploited in downstream tasks like KIE.
@@
Another category of pre-training objectives stems from the generative KIE methods, where corresponding models need to learn both natural language understanding and natural language generation (i.e., generating output sequences conditioned on input sequences) capabilities. To this end, (Cao et al., 2022a, b) adopt the pre-training tasks defined by (Dong et al., 2019), namely Unidirectional LM, Bidirectional LM and Sequence-to-Sequence LM. Importantly, all pre-training tasks are considered simultaneously by using a shared transformer network which can alternate between the three objectives. (Tang et al., 2023) propose a range of self-supervised and supervised generative pre-training objectives based on task prompts and target outputs in textual form, for example where the model must predicts missing texts and locate them in document images as a structured target sequence. The approach by (Davis et al., 2023) follows a specific strategy for learning text recognition, document understanding and generative capabilities. It generally stands out in terms of its pre-training setup, as over 25 different objectives across different document types are being used. Nonetheless, a strong focus is placed on MLM-related tasks. One important pre-training task revolves around the model predicting a structured JSON output for a given document, which is also ultimately used for KIE. This strategy can be helpful in real-world settings, in which a detailed hierarchical output of input documents is required, such as an adequate differentiation between individual line items in invoices. Non-generative KIE methods that for example decode outputs with a sequence labeling layer usually have difficulties in reconstructing such hierarchies and require additional post-processing steps. To let the model learn text recognition, (Kim et al., 2022; Dhouib et al., 2023) use a pre-training task where the model must predict the next token while considering the previous tokens as well as the document image.
@@ 6.2. Practical perspective
One work that stands out in terms of its consideration of domain knowledge is the approach by (Arroyo et al., 2022). The authors make use of such knowledge by proposing a hybrid KIE system based on both DL-models and rule-based methods including several post-processing steps to improve the extraction results. This includes, for example, the correction of automatically extracted product codes and product prices. However, this post-processing is considered as a supplement to the DL model in case of incorrect predictions and not an architecture-based adaptation. Another example is the work by (Palm et al., 2019), which integrates domain-oriented constraints regarding invoices that the total amount is the sum of the subtotal and tax total values, and that the tax total can be computed as the product of subtotal and tax percentage. The authors achieve this by adjusting the loss calculations during training, but report that no significant performance improvements were achieved. At the same time, this emphasizes the need for further research in this area (see also chapter 7). While not directly implementing such aspects in the proposed methods, some authors also designed their approaches in a way to allow for integrating domain knowledge. For example, in the work of (Tata et al., 2021), candidates are first identified for fields to be extracted and subsequently scored. In this regard, one example the authors mention is to define a scoring function that integrate constraints such as the fact that an invoice date must precede its due date.
As discussed in section 2.1, adopting a practical perspective and considering domain knowledge can provide valuable insights for an automated extraction system in various ways. However, this aspect has not been properly explored in the literature, even though it has been named as a key challenge for DU in the past (Motahari et al., 2021). Therefore, future work should consider adopting a business process perspective in order to develop approaches that explicitly exploit this unrealized potential. This could also lead to the identification of completely new potentials of KIE systems that are not yet considered in current research, as the current focus often lies only on increasing the performance with respect to established benchmark datasets. These considerations may require a deeper understanding of the associated business processes, which can be achieved by modeling computer-integrated systems including data, behavior and control flows (Fettke and Reisig, 2021). The potential of context-aware DL systems has been investigated in other research areas. For example, (Weinzierl et al., 2019) propose the fusion of information extracted from business documents with traditional event log data in order to enhance predictive process monitoring capabilities. However, the other direction, i.e., feeding context-aware data obtained from process mining techniques into KIE systems, has not yet been considered.
@@ 6.3. Trends
One can observe several trends in KIE research since 2017 in a wide variety of facets, some of which were also addressed in chapter 5. One aspect is the dominance of the three main paradigms of KIE systems and most notably sequence-based methods, which represent the predominant approach since 2021. To the contrary, grid-based methods have never been able to establish themselves, despite their comparability with graph-based approaches, which are even more popular now (especially in 2023) than they were in the beginning. At the same time, this indicates that VRDs with usually complex layouts cannot be appropriately represented as grids with well-defined structures that are less flexible compared to dynamic graphs. Also, using conjunction of different paradigms (e.g., sequence-based and graph-based) is not as widely adopted as the focus on one particular paradigm.
@@
Since 2022, an increased focus has been placed on creating OCR-independent and/or generative KIE approaches. The latter may be the consequence of the increasing popularity of generative AI methods such as LLMs in recent years and their influence on the DU domain. The increasing shift towards OCR-independent systems can be explained by the otherwise necessary external OCR engines, which can be error-prone and therefore lead to falsely extracted information, especially in real-world use cases (Cui et al., 2021).
@@
It can also be observed that over time, there have been different strategies on how to develop and improve KIE systems. While towards the beginning, a lot of research effort has been placed on how to properly integrate the different input modalities, other strategies have also emerged in recent years. For example, some publications highlight and investigate the importance of the reading order. In traditional methods, and also as a consequence of the reliance on external OCR engines, a simple top-bottom and left-right reading order is usually adopted. This however can lead to a non-ideal segmentation of complex VRDs. To this end, approaches such as those by (Peng et al., 2022; Li et al., 2023b; Shi et al., 2023; Zhang et al., 2023b) investigate more sophisticated methods to obtain an optimized reading order that better suits the actual document layout. However, what was not or hardly ever considered as a lever for better KIE performance, is the model size in terms of trainable parameters. Between 2017 and 2023, there was no significant increase in model sizes, as also discussed in connection with figure 5.
@@
It is also positive to note that code and/or model weights are being published more frequently in recent years. Even if the absolute number of implementations that have been shared is relatively low, it has become increasingly more frequent, especially since 2021. This is a positive development, as it can accelerate research progress and research dissemination. This observation also goes hand in hand with the increased popularity of HuggingFace [^8] https://huggingface.co/, which is a platform that provides tools, datasets, and pre-trained models to facilitate research in NLP and CV. Some previously discussed KIE approaches are also available, for example LayoutLMv3999https://huggingface.co/microsoft/layoutlmv3-base.
@@
It can be assumed that these trends and observations will continue in future KIE research. Especially a shift towards generative KIE systems seems evident, as corresponding models become more and more popular across multiple domains. Initial approaches that make use of powerful LLMs exist, however they do not consistently achieve competitive results compared to specialized DU methods (such as LayoutLM) yet. In this regard, additional research is required on how to close this performance gap and thus make corresponding approaches more viable, especially with regard to the trade-off between model size (and therefore hardware requirements) and extraction performance as discussed in section 5.3.2.

### v2

@@ 5.3.2. Quantitative comparison
To ensure a meaningful and representative quantitative comparison, in the following we focus on the by far most commonly used datasets CORD, FUNSD and SROIE. We also only consider the F1 score, as it is the most commonly used evaluation metric. These choices allow for the largest possible sample size and a comparability across a wider range of approaches. Note that we do not distinguish between micro, macro or weighted F1 averages [^6] For an explanation of these aggregated metrics, we refer to (Sokolova and Lapalme, 2009)., since in many cases it has not been explicitly specified by the authors. The results should therefore be treated with caution.
To this end, figure 5(c) shows the results for the three datasets given all papers that present results on the corresponding datasets. A trend line is also displayed. Only FUNSD shows an improvement over time, while for the other two datasets the trend line shows a decline. However, all graphs show a few outliers that performed worse than previous and subsequent work. One reason for this could be a different calculation of the F1 scores, e.g., by using macro averages instead of micro averages. Another possible factor could be the alignment of the models with the specific nature of the task. For example, models that do not take into account the spatial and semantic features of the datasets may perform poorly despite having a large number of parameters. These outliers will therefore negatively affect the trend line. The best performing approach for CORD is proposed by (Gu et al., 2021), for FUNSD by (Xu et al., 2024), and for SROIE by (Wang et al., 2024a). On average, very comparable results are obtained for CORD and SROIE, with an average F1 score of 0.94 and 0.95 respectively. This can be attributed to the fact that both datasets consist mainly of structured receipts, where layout and spatial encoding play a crucial role in performance. In contrast, FUNSD shows significantly worse results, with an average F1 score of only 0.83. One problem that KIE approaches face with FUNSD may be that this dataset was not primarily constructed for the KIE task. The fields to be extracted often span multiple lines of text, which is very different from traditional KIE datasets that aim at extracting key information such as a specific date.
@@
Figure 5(c) also shows the relationship between parameter count and performance across the three datasets, with the x-axis in logarithmic scale. In all cases, there is a clear clustering with some outliers. The correlation coefficients are -0.02 (p=0.90) for CORD, 0.08 (p=0.60) for FUNSD, and 0.11 (p=0.63) for SROIE, indicating no significant correlation between model size and performance. Notably, the slight negative correlation for CORD suggests that smaller models can outperform larger ones. The figure also distinguishes KIE paradigms by color. Notably, graph-based methods often appear as outliers. Especially in case of FUNSD, this observation can be attributed to its unique layout, where entities often span multiple lines, making graph representations suboptimal. Hybrid models combining graph and sequence-based approaches tend to outperform pure graph-based models, especially in this scenario. The analysis further shows that increasing model size beyond a certain threshold yields diminishing returns. This suggests that architecture design, tailored to the task, is more critical than model size. This insight is particularly useful in resource-constrained environments, where smaller, task-aligned models can achieve competitive or superior results. The performance spread in FUNSD is wider compared to CORD and SROIE, with some smaller models struggling, likely due to the complex layout and semantic requirements of the dataset. Interestingly, the LLM-based approaches often do not achieve competitive results and in some cases are among the worst performing KIE systems. Only the LLM-based approach by (Fujitake, 2024) achieves almost state-of-the-art results for CORD and FUNSD. This highlights again that larger models don’t always lead to better performance if the architecture is not suitable. On the contrary, the approach by (Gbada et al., 2024b) with only 0.8 million parameters achieves promising results and is even one of the best systems in case of CORD.
In conclusion, this analysis emphasizes that model design and task alignment are more important than model size. For structured documents as seen in CORD and SROIE, simpler models with strong layout awareness are effective, while hybrid models combining sequence and graph-based features work best for more complex KIE settings like FUNSD. Architectural choices need to balance semantic complexity with layout structure to optimize KIE performance. It is important to note that the findings are based on datasets with strong domain overlap – in particular CORD and SROIE, which consist of receipts – and that all three analyzed datasets are in English. As a result, the findings may have limited generalizability to other application domains or languages. The outlined performance trade-offs may be more pronounced for other document types and layouts, where larger models could yield significantly better results.
@@ 6. Discussion
6.1. Technical analysis
KIE approaches vary widely in how they represent and process VRDs, with strengths and weaknesses in handling multi-modal inputs, spatial document structures, and semantic richness. Sequence-based models, such as LayoutLM and its variants, dominate due to their ability to leverage large-scale pre-training and various attention mechanisms. These models incorporate spatial information via positional embeddings and often use linear classification layers to perform NER. However, such classification layers are limited in modeling label dependencies across sequences – unlike CRFs – and struggle with discontinuous entities, which are common in real-world documents (Tu et al., 2024). Moreover, while end-to-end models provide a unified extraction pipeline, OCR-dependent systems still achieve superior performance in benchmarks since they decouple the complex tasks of text recognition and entity extraction (Scius-Bertrand et al., 2024). Graph-based models excel at capturing document structure through spatially and semantically defined graph structures. At the same time, GCNs and GATs allow for effective layout-aware propagation. However, many approaches rely on heuristic graph construction, which can introduce noise or bias. In contrast, learned graph structures can show improved robustness and generalization at the expense of model complexity (Xu et al., 2024). Edge features remain underutilized, despite evidence of their potential to improve spatial reasoning. Grid-based approaches treat documents as grids and apply segmentation-like techniques for KIE. These methods are particularly suited to highly structured VRDs, but are less flexible in adapting to different layouts. They also tend to use limited semantic encoding, with little advancements beyond early design choices such as those proposed by (Katti et al., 2018). This is also reflected in the fact that research on grid-based approaches was mostly conducted before 2022. Novel LLM-based methods introduce a generative and instruction-driven paradigm for KIE. While these models offer strong generalization and task flexibility through their use of large foundation models, they have practical limitations. Layout information is often injected via simple prompt engineering, which is not native to the LLM training procedure and can impact efficiency. Furthermore, inference times are significantly higher - under certain circumstances up to 120 times higher than other approaches on comparable hardware (Perot et al., 2024) — which poses scalability challenges. Although hallucinations (i.e., the generation of text not present in the original document) are rare, occurring less than 1% in some cases (Perot et al., 2024), dedicated mechanisms may still be required to prevent them.
Almost all KIE approaches have limited capabilities to handle long and multi-page VRDs. This is mainly due to architectural constraints, e.g., transformers have a fixed input sequence length, usually 512 tokens, which is often not even enough to capture even a single page of a VRD. Graph-based and grid-based methods are unable to capture multi-page dependencies because their document representations are defined page by page. LLM-based approaches can overcome this problem to some extent, as the backbone LLMs usually provide longer context lengths (e.g., 4,096 tokens). This aspect is less of a problem when processing VRDs with fewer words (e.g., receipts as seen in CORD and SROIE), but can play a role in documents with more content, such as forms (see FUNSD). All in all, an adequate understanding of long VRDs and maintaining context over long sequences remains a challenge.
Overall, the different categories require trade-offs between semantic expressiveness, spatial reasoning, and practical constraints. Transformers offer strong contextual understanding but need improvements in sequence modeling and layout perception. Graph-based methods capture document structure explicitly and can benefit from learning the graph connectivity, rather than relying on heuristic algorithms. Grid-based models handle layout well, but lack semantic flexibility. LLM-based approaches allow for a broad adaptability, but remain computationally intensive and less effective at precise spatial reasoning.
@@ 6.2. Pre-Training and Fine-Tuning procedures
As discussed in chapter 5.2.1, the majority of authors adopt MVLM during pre-training. In some cases, it is even the only pre-training objective that is being used. Given the general popularity of MLM in the NLP domain, it is not surprising that the MVLM task, which is derived from it, is also popular in KIE research. It represents an effective method to let models learn useful representations given the surrounding textual and positional contexts. As the adequate processing of VRDs also requires to consider visual cues, newer pre-training objectives have been proposed in order to fuse visual information into KIE systems. This can be done in different ways, for example by letting the model reconstruct local or global document image snippets (Appalaraju et al., 2021; Tu et al., 2023) or predicting the segment length of an image snippet (Li et al., 2021c). Other possibilities aim at various matching-tasks between text and document images, for example where the model has to predict whether a sentence describes a document image (Appalaraju et al., 2021), if a given image and text are part of the same page (Xu et al., 2021) or if image patches of a word are masked or not (Huang et al., 2022). Some authors also define additional pre-training objectives to better exploit positional information, for example by the means of classification problems to estimate a tokens placement into a document image area (Li et al., 2021a; Wang et al., 2022b) or to estimate relative positional directions of an element to its neighbors (Li et al., 2021c). Research is also conducted related to a better understanding of numerical values and their relationships in business documents (Douzon et al., 2022). This can be beneficial for processing document types such as invoices, which usually contain various monetary values that are closely related to each other. To conclude, the proposed pre-training objectives are carefully designed to allow for an adequate fusion of textual, positional and visual inputs, which helps corresponding models to obtain a document understanding which can then be exploited in downstream tasks like KIE.
@@
Another category of pre-training objectives stems from the generative KIE methods, where corresponding models need to learn both natural language understanding and natural language generation (i.e., generating output sequences conditioned on input sequences) capabilities. To this end, (Cao et al., 2022a, b) adopt the pre-training tasks defined by (Dong et al., 2019), namely Unidirectional LM, Bidirectional LM and Sequence-to-Sequence LM. Importantly, all pre-training tasks are considered simultaneously by using a shared transformer network which can alternate between the three objectives. (Tu et al., 2023) propose a range of self-supervised and supervised generative pre-training objectives based on task prompts and target outputs in textual form, for example where the model must predicts missing texts and locate them in document images as a structured target sequence. The approach by (Davis et al., 2023) follows a specific strategy for learning text recognition, document understanding and generative capabilities. It generally stands out in terms of its pre-training setup, as over 25 different objectives across different document types are being used. Nonetheless, a strong focus is placed on MLM-related tasks. One important pre-training task revolves around the model predicting a structured JSON output for a given document, which is also ultimately used for KIE. This strategy can be helpful in real-world settings, in which a detailed hierarchical output of input documents is required, such as an adequate differentiation between individual line items in invoices. Non-generative KIE methods that for example decode outputs with a sequence labeling layer usually have difficulties in reconstructing such hierarchies and require additional post-processing steps. To let the model learn text recognition, (Kim et al., 2022; Dhouib et al., 2023) use a pre-training task where the model must predict the next token while considering the previous tokens as well as the document image.
@@ 6.3. Practical perspective
One work that stands out in terms of its consideration of domain knowledge is the approach by (Arroyo et al., 2022). The authors make use of such knowledge by proposing a hybrid KIE system based on both DL-models and rule-based methods including several post-processing steps to improve the extraction results. This includes, for example, the correction of automatically extracted product codes and product prices. However, this post-processing is considered as a supplement to the DL model in case of incorrect predictions and not an architecture-based adaptation. Another example is the work by (Palm et al., 2019), which integrates domain-oriented constraints regarding invoices that the total amount is the sum of the subtotal and tax total values, and that the tax total can be computed as the product of subtotal and tax percentage. The authors achieve this by adjusting the loss calculations during training, but report that no significant performance improvements were achieved. At the same time, this emphasizes the need for further research in this area (see also chapter 7). While not directly implementing such aspects in the proposed methods, some authors also designed their approaches in a way to allow for integrating domain knowledge. For example, in the work of (Tata et al., 2021), candidates are first identified for fields to be extracted and subsequently scored. In this regard, one example the authors mention is to define a scoring function that integrates constraints such as the fact that an invoice date must precede its due date.
As discussed in section 2.1, adopting a practical perspective and considering domain knowledge can provide valuable insights for an automated extraction system in various ways. However, this aspect has not been properly explored in the literature, even though it has been named as a key challenge for DU in the past (Motahari et al., 2021). Therefore, future work should consider adopting a business process perspective in order to develop approaches that exploit this unrealized potential. This could also lead to the identification of completely new potentials of KIE systems that are not yet considered in current research, as the current focus often lies only on increasing the performance with respect to established benchmark datasets. These considerations may require a deeper understanding of the associated business processes, which can be achieved by modeling computer-integrated systems including data, behavior and control flows (Fettke and Reisig, 2021). The potential of context-aware DL systems has been investigated in other research areas. For example, (Weinzierl et al., 2019) propose the fusion of information extracted from business documents with traditional event log data in order to enhance predictive process monitoring capabilities. However, the other direction, i.e., feeding context-aware data obtained from process mining techniques into KIE systems, has not yet been considered.
@@ 6.4. Trends
One can observe several trends in KIE research since 2017 in a wide variety of facets. One aspect is the dominance of certain paradigms of KIE systems – most notably sequence-based methods, which represent the predominant approach since 2021. To the contrary, grid-based methods have never been able to establish themselves, despite their comparability with graph-based approaches, which have been popular since the beginning. At the same time, this indicates that VRDs with usually complex layouts cannot be appropriately represented as grids with well-defined structures that are less flexible compared to dynamic graphs. As expected, LLM-based systems have emerged in 2023 and quickly became popular. Also, using conjunctions of different paradigms (e.g., sequence-based and graph-based) is not as widely adopted as the focus on one particular paradigm.
@@
Since 2022, an increased focus has been placed on creating OCR-independent and generative KIE approaches. The latter is the consequence of the increasing popularity of generative AI methods such as LLMs in recent years and their influence on the DU domain. The increasing shift towards OCR-independent systems can be explained by the otherwise necessary external OCR engines, which can be error-prone and therefore lead to falsely extracted information, especially in real-world use cases (Cui et al., 2021).
@@
It can also be observed that over time, there have been different strategies on how to develop and improve KIE systems. While towards the beginning, a lot of research effort has been placed on how to properly integrate the different input modalities, other strategies have also emerged in recent years. For example, some publications highlight and investigate the importance of the reading order. In traditional methods, also as a consequence of the reliance on external OCR engines, a simple top-bottom and left-right reading order is usually adopted. This however can lead to a suboptimal segmentation of complex VRDs. To this end, approaches such as those by (Peng et al., 2022; Tang et al., 2023; Zhang et al., 2023b) investigate more sophisticated methods to obtain an optimized reading order that better suits the actual document layout. (Villanova-Aparisi et al., 2024) also propose evaluation metrics that overcome biases resulting from the reading order. However, what was not or hardly ever considered as a lever for better KIE performance, is the model size in terms of trainable parameters. Between 2017 and 2024, there was no significant increase in model sizes besides a more frequent use of LLMs.
@@
It is also positive to note that code and/or model weights are being published more frequently in recent years. Even if the absolute number of implementations that have been shared is relatively low, it has become increasingly more frequent, especially since 2021. This is a positive development, as it can accelerate research progress and research dissemination. This observation also goes hand in hand with the increased popularity of HuggingFace [^7] https://huggingface.co/, which is a platform that provides tools, datasets, and pre-trained models to facilitate research in NLP and CV. Some previously discussed KIE approaches are also available, for example LayoutLMv3888https://huggingface.co/microsoft/layoutlmv3-base.
@@
It can be assumed that these trends and observations will continue in future KIE research. Especially a shift towards generative KIE systems seems evident, as corresponding models become more and more popular across multiple domains. Initial approaches that make use of powerful LLMs exist, however they do not consistently achieve competitive results compared to specialized DU methods yet. In this regard, additional research is required on how to close this performance gap and thus make corresponding approaches more viable, especially with regard to the trade-off between model size (and therefore hardware requirements) and extraction performance as discussed in section 5.3.2.